# Harsha Kokel IBM Research

### v3

@@ I Introduction
  • Demonstration of TabSketchFM on search. We compared the performance of TabSketchFM’s fine tuned models against neural and non-neural baselines. In the case of join search, TabSketchFM outperformed the nearest systems geared for join search by almost 70% in F1 scores, but tellingly, the sentence encoding model outperformed existing systems as well by 63%. By combining the two embeddings (from TabSketchFM and sentence encodings), we gain almost 73% over other systems in F1.
@@ III-A Sketches Preparation
MinHash Sketches. For each column, we treat cell-value as a string and compute a MinHash signature from the set of cell-values.555 While it may not make sense to convert float or integer columns into MinHash signatures, we do so because it is often difficult to tell if a column is truly a float or an integer or is really a categorical value. For string columns, we also compute a MinHash signature for set of words within the column. The rationale for this second MinHash is that the words can sometimes capture its semantics (e.g., if street appears in two different columns, they maybe both about address, which is useful if one considers a union of two tables that do not necessarily share cell values). Consider the Reference Area column in the table shown in Fig. 1. The cell-values MinHash Sketch treats "Austria Vienna" as an element in the set of strings, whereas "Austria" and "Vienna" are treated as separate strings in words-based MinHash signature. Both MinHashes are concatenated into a single input vector for string columns. For numerical and date columns, only the MinHash for the cell values is included in the input vector.
@@ III-B Input Representation
2. Token Position Embeddings. Considering each column analogous to a sentence in a paragraph, we re-purpose token positional embeddings to reflect a token’s position within a column name. For instance, as shown in Fig. 1 with light gray boxes, the token position of the token Area is 1 because it is the second token in that column (after Reference).
3. Column Position Embeddings. Columns themselves have positions and are encoded through a new embedding layer which can range from 1 to the total number of columns, [^6] Position 0 is reserved for any tokens in the table description. as shown in green boxes in Fig. 1. The rationale for including column positions is that they sometimes do have semantics; e.g., a street address next to the city, followed by a zip code. Of course, table semantics do not change as a function of column order; nevertheless, we included a position in case it helps the model understand column semantics, with the assumption that the attentional mechanism would disregard order if necessary.
4. Column Type Embeddings. The data types of columns: string, date, integer, or float are encoded through additional embedding. In order to determine column type, we made a best-case effort to parse the first 10 values of each column as dates, integers, or floats and defaulted to string if we could not convert them. In mixed-type columns, this can yield poor results, but at the very worst, at least one of the types was assigned to these columns. This is shown in Fig. 1 in orange, where string, integer, float and date types are represented with integers 1, 2, 3, and 4, resp.
@@
Figure 2: (a) For pretraining, the table input embedding (see Section III-B) is passed through 12-layered best encoder to generate output token embeddings. The output token embeddings are used for MLM objective with cross entropy loss( Section III-C). We initialize the weight of the BERT encoder as well as the token embeddings from the pretrained base uncased BERT model. (b) Two input tables are concatenated and passed through the pretrained TabSketchFM. The BERT pooler output (embedding of the first token) is taken and passed through a dropout and a linear layer to generate output of size N.
@@ III-C Pretraining
Data Augmentation. In the image processing literature [53], different transformations are applied over an original image to generate new image samples for training. This makes the trained model robust on different variants of the same image. We want the model to be robust to the order of columns such that shuffling columns in a table does not impact its semantics. So, we created three different versions of the table, by changing the column order, which in turn, changed the table’s content snapshot (see Section III). This increases the total number of pretraining tables to 290,948 and for each of them, we create signatures as explained in Section III.
@@
$\frac{-1}{N}(\sum_{n=1}^{N}y_{i}*log(\hat{y_{i}}))$ (1)
Equation 1 defines cross-entropy loss, where $y_{i}$ is the class labels in the ground truth, $\hat{y_{i}}$ is the predicted class label and N is the number of training examples in the train set.
@@
In our scenario, for each table, we mask a single column name such that all tokens corresponding to that column are masked. This is analogous to the natural language literature’s whole word masking where all tokens corresponding to a word are masked. In addition, we used MLM probability to mask the tokens in the table description as well. As shown in Fig. 3, we generate several samples from a single table. Note that each masking gives us an example. For small tables with less than $5$ columns, we mask each of its columns one after another. However, if a table has a large number of columns, this produces a lot of examples for the same table, over-representing it. So, for large tables with more than $5$ columns, we randomly select five columns and mask them. Following this strategy over the augmented pretraining dataset, we get $730,553$ examples in training, $54,430$ examples in validation, and $58,141$ examples in test sets. We provide the experimental setup for pretraining in Section IV. Figure 2(a) shows the TabSketchFM pretraining architecture.
@@ IV Experiments
ECB Union (R2) $0.03\pm 0.03$ $-0.01\pm 0.01$ $0.02\pm 0.01$ $\underline{0.87\pm 0.01}$ $0.35\pm 0.06$ $\boldsymbol{0.90\pm 0.03}$
Wiki Jaccard (R2) $00.00\pm 0$ 0.00 $-0.03\pm 0.03$ $0.25\pm 0.02$ $\underline{0.43\pm 0.02}$ $0.33\pm 0.03$ $\boldsymbol{0.58\pm 0.03}$
Wiki Containment (R2) $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.21\pm 0.01$ $\underline{0.35\pm 0.03}$ $0.30\pm 0.03$ $\boldsymbol{0.58\pm 0.02}$
@@
We pretrain TabSketchFM using $4$ A100 $40$GB GPUs for $2$ days, until the model converged. [^8] The source code is available on https://github.com/ibm/tabsketchfm Our pretrained model contains $118M$ parameters, similar to other row-based models in our experiments. We use patience of $5$, i.e., we consider the model as having converged if the validation loss does not decrease for more than $5$ epochs. As studied in the literature [58] this not only reduces training time but also helps us avoid overfitting. For finetuning the cross encoders, for all tasks, we use one A100 $40$GB GPU. Most finetuning tasks converged within $6$ hours using the same 5 epoch patience as pretraining except for the Wiki Union benchmark which took 24 hours. Our code is accessible at https://anonymous.4open.science/r/tabsketchfm.
@@ IV-A2 Results
Table II compares the performance of TabSketchFM with baseline models on LakeBench. For regression tasks, we report R2 statistics, and for (binary and multiclass) classification tasks, we report a weighted F1 score to handle skew in classes [^10] Our metrics were implemented from scikit-learn: https://scikit-learn.org/stable/modules/model_evaluation.html. The best performance is highlighted in bold and the second best is underlined. All the results presented are aggregated over 5 random seeds. We see that on five of the eight tasks, TabSketchFM outperformed all baselines, performed as well as the best methods on one, and performed second best on the remaining two.
@@
We also performed an error analysis on table pairs that TabSketchFM got wrong on Wiki-Union, Spider open data and CKAN. In all these cases, mean hamming distances of MinHash sketches for positive pairs were equivalent to those from negative pairs; suggesting that the sketches alone did not contain sufficient information to discriminate those examples. TaBERT’s superior performance on the Wiki Union case, combined with the fact that lack of discrimination of MinHashes provides a clue to at least one weakness in the current version of TabSketchFM, which is that we do not consider any values from the column at all. As shown in Figure 5, two columns may be unionable when their values map to the same domain (e.g., different municipalities in Slovakia), with minimal overlap in actual values. In such cases, a system such as TaBERT can still determine that the cell values are similar, and belong to the same semantic type, whereas TabSketchFM cannot. We show how to incorporate cell values externally in the section on search which addresses this issue.
@@ IV-B Effectiveness of sketch types on task types
To explore the relevance of the sketches and content snapshots for the different data discovery tasks, we conduct an ablation study where we use only one type of sketch. The results on LakeBench tasks with the random seed 0 are shown in Table III. In  Table IV we present another ablation study where we remove one sketch and retained the rest. Not surprisingly, MinHash sketches are crucial for join tasks, where MinHash alone seems to dominate overall performance of TabSketchFM. Subset selection clearly relies on the overlap of the data distribution; numerical sketches alone match the performance of TabSketchFM. Given that almost 69% of the columns were numerical in the subset benchmark, this outcome might be expected. The effect of content snapshot was less clear. Looking at Table IV, it is evident that the removal of content snapshot does in fact contribute somewhat to the Wiki join tasks. Numerical sketches contribute to ECB union which is perhaps because approximately 50% of the columns are not strings. Removal of MinHashes mirrors the result from earlier, which is, join tasks are most affected. Surprisingly, the removal of MinHashes affected the subset task; even though numerical sketches by themselves produced good performance. It is possible that there are complex interactions between the sketches. In summary, we answer Q2 from our ablation studies that different sketches play different roles across tasks, and may even interact to determining data discovery performance.
@@ IV-C Application to search
TaBERT-FT 5.88 0.43 0.04
LSH-Forest 10.48 0.8 0.08
Josie 19.56 0.99 0.12
DeepJoin 18.88 0.96 0.11
WarpGate 18.58 0.95 0.11
SBERT 83.67 0.95 0.89
@@
TABLE VII: F1, Precision & Recall for TUS union search.
@@
TaBERT-FT 26.66 0.90 0.30
TUTA-FT 27.27 0.89 0.31
Starmie 27.48 0.96 0.32
D3L 18.98 0.75 0.21
SANTOS 20.83 0.81 0.23
SBERT 31.13 0.99 0.36
TabSketchFM 30.43 0.97 0.35
TabSketchFM-SBERT 30.72 0.99 0.35
@@ IV-C1 Join Search
Figure 4(a) shows the significant advantage of TabSketchFM over all other baselines, and furthermore, adding column embeddings using SBERT improved performance for joins. Table IV-C shows the mean F1 scores for all methods. In this and other benchmarks, the best performing system and the runner up were better than all other methods, using statistical significance tests. Table IV-C presents the Mean F1, Precision @ 10, and Recall @10 values.
@@ IV-C2 Union Search
    For each column $c$ in a given table $t$, KnnSearch returns the nearest $k*3$ columns in the dataset. Note that we try to get a lot more columns than $k$ simply to increase the number of candidate columns we consider at this step because multiple columns from a single table might match a given column.
@@
For each technique, we use default parameters suggested in the respective papers. For TUTA we had to use table embeddings because that was what the model made available. For both the TUTA and TaBERT, we use the model that was fine tuned on TUS-SANTOS dataset from LakeBench in Sec. IV-A(denoted by TUTA-FT and TaBERT-FT). The results for TabSketchFM compared to the baselines show that TabSketchFM matched the best system [16] on SANTOS, when we added embeddings from SBERT, see Figure 4(b), and Table IV-C for a summary across methods. For TUS, value embeddings from the off-the-shelf SBERT model performed very well, as did TabSketchFM when enhanced with value embeddings; see Figure 4(c) and Table IV-C. This suggests for union, column values alone are sufficient to find good unionable results; table specific embeddings do not seem to add much. Note that value embeddings out performed all the other systems, pointing to the power of existing pretrained models in embedding tabular data.

### v4

@@ I Introduction
  • Demonstration of TabSketchFM on search. We compared the performance of TabSketchFM’s fine tuned models against neural and non-neural baselines. In the case of join search, TabSketchFM outperformed most other systems geared for join search in terms of F1 scores.
@@ III-A Sketches Preparation
MinHash Sketches. For each column, we treat cell-value as a string and compute a MinHash signature from the set of cell-values.555 While it may not make sense to convert float or integer columns into MinHash signatures, we do so because it is often difficult to tell if a column is truly a float or an integer or is really a categorical value. For string columns, we also compute a MinHash signature for set of words within the column. The rationale for this second MinHash is that the words can sometimes capture its semantics (e.g., if street appears in two different columns, they maybe both about address, which is useful if one considers a union of two tables that do not necessarily share cell values). Consider the Reference Area column in the table shown in Fig.˜1. The cell-values MinHash Sketch treats "Austria Vienna" as an element in the set of strings, whereas "Austria" and "Vienna" are treated as separate strings in words-based MinHash signature. Both MinHashes are concatenated into a single input vector for string columns. For numerical and date columns, only the MinHash for the cell values is included in the input vector.
@@ III-B Input Representation
2. Token Position Embeddings. Considering each column analogous to a sentence in a paragraph, we re-purpose token positional embeddings to reflect a token’s position within a column name. For instance, as shown in Fig.˜1 with light gray boxes, the token position of the token Area is 1 because it is the second token in that column (after Reference).
3. Column Position Embeddings. Columns themselves have positions and are encoded through a new embedding layer which can range from 1 to the total number of columns, [^6] Position 0 is reserved for any tokens in the table description. as shown in green boxes in Fig.˜1. The rationale for including column positions is that they sometimes do have semantics; e.g., a street address next to the city, followed by a zip code. Of course, table semantics do not change as a function of column order; nevertheless, we included a position in case it helps the model understand column semantics, with the assumption that the attentional mechanism would disregard order if necessary.
4. Column Type Embeddings. The data types of columns: string, date, integer, or float are encoded through additional embedding. In order to determine column type, we made a best-case effort to parse the first 10 values of each column as dates, integers, or floats and defaulted to string if we could not convert them. In mixed-type columns, this can yield poor results, but at the very worst, at least one of the types was assigned to these columns. This is shown in Fig.˜1 in orange, where string, integer, float and date types are represented with integers 1, 2, 3, and 4, resp.
@@
Figure 2: (a) For pretraining, the table input embedding (see Section˜III-B) is passed through 12-layered best encoder to generate output token embeddings. The output token embeddings are used for MLM objective with cross entropy loss( Section˜III-C). We initialize the weight of the BERT encoder as well as the token embeddings from the pretrained base uncased BERT model. (b) Two input tables are concatenated and passed through the pretrained TabSketchFM. The BERT pooler output (embedding of the first token) is taken and passed through a dropout and a linear layer to generate output of size N.
@@ III-C Pretraining
Data Augmentation. In the image processing literature [53], different transformations are applied over an original image to generate new image samples for training. This makes the trained model robust on different variants of the same image. We want the model to be robust to the order of columns such that shuffling columns in a table does not impact its semantics. So, we created three different versions of the table, by changing the column order, which in turn, changed the table’s content snapshot (see Section˜III). This increases the total number of pretraining tables to 290,948 and for each of them, we create signatures as explained in Section˜III.
@@
$)\frac{-1}{N}(\sum_{n=1}^{N}y_{i}*log(\hat{y_{i}}$)) (1)
Equation˜1 defines cross-entropy loss, where $y_{i}$ is the class labels in the ground truth, $\hat{y_{i}}$ is the predicted class label and N is the number of training examples in the train set.
@@
In our scenario, for each table, we mask a single column name such that all tokens corresponding to that column are masked. This is analogous to the natural language literature’s whole word masking where all tokens corresponding to a word are masked. In addition, we used MLM probability to mask the tokens in the table description as well. As shown in Fig.˜3, we generate several samples from a single table. Note that each masking gives us an example. For small tables with less than $5$ columns, we mask each of its columns one after another. However, if a table has a large number of columns, this produces a lot of examples for the same table, over-representing it. So, for large tables with more than $5$ columns, we randomly select five columns and mask them. Following this strategy over the augmented pretraining dataset, we get $730,553$ examples in training, $54,430$ examples in validation, and $58,141$ examples in test sets. We provide the experimental setup for pretraining in Section˜IV. Figure 2(a) shows the TabSketchFM pretraining architecture.
@@ IV Experiments
ECB Union (R2) $03\pm 0.03 0.0$$1\pm 0.01 0.0$$02\pm 0.01$ $\underline{0.87\pm 0.01}$ $0.35\pm 0.06$ $\boldsymbol{0.90\pm 0.03}$
Wiki Jaccard (R2) $00\pm 0.00 0.0$$03\pm 0.03$ $0.25\pm 0.02$ $\underline{0.43\pm 0.02}$ $0.33\pm 0.03$ $\boldsymbol{0.58\pm 0.03}$
Wiki Containment (R2) $\pm 0.00 0.00±0.00$$\pm 0.00$ $0.21\pm 0.01$ $\underline{0.35\pm 0.03}$ $0.30\pm 0.03$ $\boldsymbol{0.58\pm 0.02}$
@@
We pretrain TabSketchFM using $4$ A100 4040GB GPUs for $2$ days, until the model converged. [^8] The source code is available on https://github.com/ibm/tabsketchfm Our pretrained model contains $118M$ parameters, similar to other row-based models in our experiments. We use patience of $5$, i.e., we consider the model as having converged if the validation loss does not decrease for more than $5$ epochs. As studied in the literature [58] this not only reduces training time but also helps us avoid overfitting. For finetuning the cross encoders, for all tasks, we use one A100 4040GB GPU. Most finetuning tasks converged within $6$ hours using the same 5 epoch patience as pretraining except for the Wiki Union benchmark which took 24 hours. Our code is accessible at https://github.com/ibm/tabsketchfm.
@@ IV-A2 Results
Table˜II compares the performance of TabSketchFM with baseline models on LakeBench. For regression tasks, we report R2 statistics, and for (binary and multiclass) classification tasks, we report a weighted F1 score to handle skew in classes [^10] Our metrics were implemented from scikit-learn: https://scikit-learn.org/stable/modules/model_evaluation.html. The best performance is highlighted in bold and the second best is underlined. All the results presented are aggregated over 5 random seeds. We see that on five of the eight tasks, TabSketchFM outperformed all baselines, performed as well as the best methods on one, and performed second best on the remaining two.
@@
We also performed an error analysis on table pairs that TabSketchFM got wrong on Wiki-Union, Spider open data and CKAN. In all these cases, mean hamming distances of MinHash sketches for positive pairs were equivalent to those from negative pairs; suggesting that the sketches alone did not contain sufficient information to discriminate those examples. TaBERT’s superior performance on the Wiki Union case, combined with the fact that lack of discrimination of MinHashes provides a clue to at least one weakness in the current version of TabSketchFM, which is that we do not consider any values from the column at all. As shown in Figure˜5, two columns may be unionable when their values map to the same domain (e.g., different municipalities in Slovakia), with minimal overlap in actual values. In such cases, a system such as TaBERT can still determine that the cell values are similar, and belong to the same semantic type, whereas TabSketchFM cannot. We show how to incorporate cell values externally in the section on search which addresses this issue.
@@ IV-B Effectiveness of sketch types on task types
To explore the relevance of the sketches and content snapshots for the different data discovery tasks, we conduct an ablation study where we use only one type of sketch. The results on LakeBench tasks with the random seed 0 are shown in Table˜III. In  Table˜IV we present another ablation study where we remove one sketch and retained the rest. Not surprisingly, MinHash sketches are crucial for join tasks, where MinHash alone seems to dominate overall performance of TabSketchFM. Subset selection clearly relies on the overlap of the data distribution; numerical sketches alone match the performance of TabSketchFM. Given that almost 69% of the columns were numerical in the subset benchmark, this outcome might be expected. The effect of content snapshot was less clear. Looking at Table˜IV, it is evident that the removal of content snapshot does in fact contribute somewhat to the Wiki join tasks. Numerical sketches contribute to ECB union which is perhaps because approximately 50% of the columns are not strings. Removal of MinHashes mirrors the result from earlier, which is, join tasks are most affected. Surprisingly, the removal of MinHashes affected the subset task; even though numerical sketches by themselves produced good performance. It is possible that there are complex interactions between the sketches. In summary, we answer Q2 from our ablation studies that different sketches play different roles across tasks, and may even interact to determining data discovery performance.
@@ IV-C Application to search
TaBERT-FT 30.16 0.43 0.32
LSH-Forest 50.84 0.8 0.7
Josie 94.86 0.99 1.00
DeepJoin 91.59 0.96 0.97
WarpGate 90.34 0.95 0.95
SBERT 83.67 0.96 0.89
@@
TABLE VII: F1, Precision & Recall for TUS union search. SANTOS results are reported only on 125 queries.
@@
TaBERT-FT 28.05 0.9 0.32
TUTA-FT 28.68 0.89 0.33
Starmie 28.79 0.9 0.33
D3L 20.77 0.6 0.23
SANTOS 24.27 0.81 0.27
SBERT 32.73 0.99 0.38
TabSketchFM 32.0 0.97 0.37
TabSketch-SBERT 32.3 0.99 0.38
@@ IV-C1 Join Search
Figure 4(a) shows an advantage of TabSketchFM over other join search baselines in Wiki Join Search benchmark. Table IV-C summarizes these results considering Mean F1, Precision @ 10, and Recall @10. In this and other benchmarks, TabSketchFM gives comparable performance with the exact match baseline Josie, and is among the top-2 best performing methods throughout the experiments. Specifically, the variation of TabSketchFM that adds SBERT column embeddings (TabSketchFM-BERT) improves performance of join search by around 3 % in terms of Mean F1.
@@ IV-C2 Union Search
    For each column $c$ in a given table $t$, KnnSearch returns the nearest k3k*3 columns in the dataset. Note that we try to get a lot more columns than $k$ simply to increase the number of candidate columns we consider at this step because multiple columns from a single table might match a given column.
@@
For each technique, we use default parameters suggested in the respective papers. For TUTA we had to use table embeddings because that was what the model made available. For both the TUTA and TaBERT, we use the model that was fine tuned on TUS-SANTOS dataset from LakeBench in Sec. IV-A(denoted by TUTA-FT and TaBERT-FT). The results for TabSketchFM compared to the baselines show that TabSketchFM matched the best system [16] on SANTOS, when we added embeddings from SBERT, see Figure 4(b), and Table IV-C for a summary across methods. For TUS, following the literaure [6, 16, 26], we report the results for tables having at least 60 unionable tables as we consider the value of k up to 60. Additionally, as SANTOS required manual annotation of intent, we only run this on a subset of $\mathbf{TUS_{Small}}$ queries used in the original paper [6]. The value embeddings from the off-the-shelf SBERT model performed very well, as did TabSketchFM when enhanced with value embeddings; see Figure 4(c) and Table IV-C. This suggests for union, column values alone are sufficient to find good unionable results; table specific embeddings do not seem to add much. Note that value embeddings out performed all the other systems, pointing to the power of existing pretrained models in embedding tabular data.